### v3

# Approximation rates for finite mixtures of location-scale models

Abstract
Finite mixture models provide useful methods for approximation and estimation of probability density functions. In the context of location-scale mixture models, we improve upon previous results to show that it is possible to quantitatively bound the ${\cal L}_{p}$ approximation error in terms of the number of mixture components $m$ for any $p\in\left(1,\infty\right)$. In particular, for approximation on $\mathbb{R}^{d}$ of targets $f_{0}\in{\cal W}^{s,p}$, the (fractional) Sobolev spaces with smoothness order $s\in\left(0,1\right]$, if $p<2$ then the error has size $O\left(m^{-\frac{s}{sq+d}}\right)$, while if $p\geq 2$, the error has size $O\left(m^{-\frac{2s}{2\left(sq+d\right)}}\right)$, where $q=1/\left(p-1\right)$. We demonstrate that these results can be used to derive estimation rates for a location-scale mixture-based adaptive least-squares estimator.
1 Introduction
1.1 Background and prior results
Finite mixture models appear ubiquitously throughout the statistics literature and their application and theory have been widely and comprehensively espoused in classic works such as Everitt & Hand, (1981), Titterington et al., (1985) McLachlan & Basford, (1988), Lindsay, (1995); McLachlan & Peel, (2000), and Frühwirth-Schnatter, (2006), along with more recent texts such as McNicholas, (2015), Ng et al., (2019), Robert et al., (2019), Chen, (2023), and Yao & Xiang, (2024), among many other sources. In application, finite mixtures are typically employed to perform one of two primary tasks: to provide a statistical model for a population that is hypothesized or known to have a heterogenous subpopulation structure, or to provide a flexible approximation to some unknown probability distribution. When the finite mixtures are employed for the latter task, the Gaussian mixture model is typically considered for its folkloric ability to provide accurate approximations of typically encountered probability models (cf. Nguyen & McLachlan,, 2019 and references therein).
On $\mathbb{R}^{d}$, the fact that linear combinations of Gaussian kernels can approximate sufficiently regular functions in ${\cal L}_{p}$ spaces goes back to the Tauberian theory of Wiener, (1932) (cf. Achieser,, 1956, Ch. VI). Subsequent generalizations and developments on such results appear throughout the approximation theory literature, and include the works of Park & Sandberg, (1991), Park & Sandberg, (1993), Xu et al., (1993), Sandberg, (2001), and Nestoridis et al., (2011)—see also Cheney & Light, (2000, Ch. 20–24).
When we restrict the approximand $f_{0}:\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ to be a probability density function, and the approximator class ($\cal A$}, say) to be convex combinations of probability density kernels, the approximation problem requires different tools. When $f_{0}$ is within the closure of $\cal A$}, and the class $\cal A$} is uniformly bounded in ${\cal L}_{p}$ norm, a Maurey–Barron–Jones type result of Donahue et al., (1997) provides an approximation theorem with known and optimal rate (see, e.g., Zeevi & Meir,, 1997 and Nguyen & McLachlan,, 2019 for example). However, when $f_{0}$ is not in the closure of $\cal A$} or when $\cal A$} is unbounded in ${\cal L}_{p}$ norm, the study of approximation demands more care.
In Bacharoglou, (2010), it is proved that by taking $\cal A$} to be the class of convex combinations of Gaussian densities on $\mathbb{R}$, one can arbitrarily well approximate $f_{0}$ in any ${\cal L}_{p}$ norm ($p\in\left[1,\infty\right]$) on $\mathbb{R}$, under integrability and tail conditions. Then, in Nguyen et al., (2020) and Nguyen et al., 2023b , generalizations are obtained, whereupon it is proved that on $\mathbb{R}^{d}$, $f_{0}$ can be approximated arbitrarily well when $\cal A$} is taken to be location-scale mixtures of some fixed probability density function, under various regularity conditions.
1.2 Main contributions
In this work, we extend upon the qualitative results of Nguyen et al., (2020) and Nguyen et al., 2023b , by proving that it is also possible to obtain the ${\cal L}_{p}$ approximation rate in the number of mixture components (i.e., number of vertices of the approximating convex combination) when conducting approximation using location-scale mixture models when some smoothness assumption on $f_{0}$ hold. Particularly, for classes of targets $f_{0}$ in the Sobolev space ${\cal W}^{1,p}\left(\mathbb{R}^{d}\right)$ (Leoni,, 2017) or the fractional Sobolev spaces ${\cal W}^{s,p}\left(\mathbb{R}^{d}\right)$ ($s\in\left(0,1\right)$; Leoni,, 2023) with bounded norms, $f_{0}$ has smoothness coefficient $\alpha=1$ or $\alpha=s$, respectively. Then, for every $f_{0}$ whose Sobolev (or fractional Sobolev) norm is bounded, if $\cal A$} and $f_{0}$ satisfy appropriate integrability assumptions, then there exist sequences of location-scale mixtures $\left(f_{m}\right)_{m}\subset{\cal A}$ such that
$\left\|f_{m}-f_{0}\right\|_{p}\apprle\begin{cases}m^{-\frac{\alpha}{\alpha q+d}}&\text{if }1<p<2\text{,}\\ m^{-\frac{\alpha q}{2\left(\alpha q+d\right)}}&\text{if }2\leq p<\infty\text{,}\end{cases}$
where $1/p+1/q=1$.
As an application of our result, we revisit the problem of adaptive least-squares density estimation considered in Klemelä, (2007) in a joint approximation and estimation setup, where the the expected squared ${\cal L}_{2}$ error between the least-squares estimator and the target $f_{0}$ is shown to converge at rate $n^{-1/2}$ when $f_{0}$ is in the closure of $\cal A$}, by taking estimator sequences with numbers of mixture components increasing at rate $n^{1/2}$ (see, also Klemela,, 2009, Ch. 19). We provide an analogous result for situations when $f_{0}$ is not in the closure of $\cal A$}, and demonstrate that convergence rate of $n^{-\alpha/\left(2\alpha+d\right)}$ can be achieved by similarly using estimators whose components number increase at the same rate.
1.3 Related literature
Before we proceed to the body of our manuscript, we note that the literature on density estimation and approximation by finite mixture models is far broader and richer than what has been mentioned thus far. An incomplete list of such works in the area are as follows. For instance, the results of Nguyen et al., (2020) and Nguyen et al., 2023b have been extended to the mixture of experts setting of Jacobs et al., (1991) in Nguyen et al., (2021) and to the spherical manifold setting in Ng & Kwong, (2022). Schemes relating to weak approximations by finite mixture models appear in Dalal & Hall, (1983), Petrone & Veronese, (2010), and Lee & Lin, (2012), while best approximation and estimation results in various divergence measures are obtained by Ma et al., (2025) for the case where $f_{0}$ is in the closure of the Gaussian mixture models. Joint approximation and estimation results, analogous to Klemelä, (2007), have been investigated by Li & Barron, (1999), Rakhlin et al., (2005), and Dalalyan & Sebbar, (2018) in the Kullback–Leibler (KL) setting; by Genovese & Wasserman, (2000) in the Hellinger distance setting; and with a modified KL divergence in Chiu Chong et al., (2024). Beyond density estimation and approximation using adaptive least-squares estimators, there also exists a rich body of literature on parameter estimation via maximum likelihood estimation for mixture models (see e.g., Nguyen,, 2013, Shen et al.,, 2013, Ho & Nguyen, 2016a, , and Ho & Nguyen, 2016b, ) and for the related mixtures of experts class (see e.g., Jiang & Tanner,, 1999, Nguyen et al., 2023a, , Nguyen et al., 2024a, , Nguyen et al., 2024b, , and Nguyen et al., 2024c, ).
The rest of the manuscript proceeds as follows. In Section 2, we outline the notational conventions of the manuscript and provide some preliminary setup. In Section 3, we prove the main approximation theoretic result, and then in Section 4, we apply the approximation theorem to obtain the previewed least-squares estimation rate. A discussion of the work is provided in Section 5. The Appendix then contains the proof of the main result and collects some required technical results involved in our proofs.
2 Preliminaries
${\cal L}_{p}$ spaces
For each $p\in\left[1,\infty\right)$, say that $f\in{\cal L}_{p}\left(\mathbb{R}^{d}\right)$ if $\left\|f\right\|_{p}<\infty$, where $\left\|f\right\|_{p}=\left(\int_{\mathbb{R}^{d}}\left|f\left(x\right)\right|^{p}\text{d}x\right)^{1/p}$ denotes the corresponding norm of the space. When $p=\infty$, we say that $f\in{\cal L}_{\infty}\left(\mathbb{R}^{d}\right)$ if $\left\|f\right\|_{\infty}=\inf\left\{c\geq 0\text{ | }\left|f\left(x\right)\right|\leq c\text{ for }\lambda\text{-a.e. }x\right\}<\infty$, where $\lambda$ always denotes the Lebesgue measure on an appropriate set. When there is no confusion, we will simplify ${\cal L}_{p}\left(\mathbb{R}^{d}\right)$ to ${\cal L}_{p}$, along with other such notation. Note that for vectors $x\in\mathbb{R}^{d}$, we will also write $\left\|x\right\|_{p}$ to denote the $\ell_{p}$ norm of the vector, unless it is misleading to do so.
Sobolev spaces
For each $p\in\left[1,\infty\right]$, we say that $f$ admits first-order weak derivatives in ${\cal L}_{p}$ if for each $k\in\left[d\right]=\left[1,d\right]\cap\mathbb{N}$, there exists a $v_{k}\in{\cal L}_{p}$, such that
$\int_{\mathbb{R}^{d}}f\left(x\right)D_{k}u\left(x\right)\text{d}x=-\int_{\mathbb{R}^{d}}v_{k}\left(x\right)u\left(x\right)\text{d}x\text{,}$
for all smooth and compactly supported functions $u\in{\cal C}_{c}^{\infty}\left(\mathbb{R}^{d}\right)$, where $D_{k}u=\partial u/\partial x_{k}$ and we write $v_{k}=D_{k}f$. The vector of first-order weak derivatives can then be written as $x\mapsto\nabla f=\left(D_{1}f,\dots,D_{d}f\right)\left(x\right)$ and we say that $f\in{\cal W}^{1,p}\left(\mathbb{R}^{d}\right)$ if
$\left\|f\right\|_{{\cal W}^{1,p}}=\left\|f\right\|_{p}+\sum_{k=1}^{d}\left\|D_{k}f\right\|_{p}<\infty\text{.}$
Similarly, for each $p\in\left[1,\infty\right)$ and $s\in\left(0,1\right)$, we say that $f\in{\cal W}^{s,p}\left(\mathbb{R}^{d}\right)$ if
$\left\|f\right\|_{{\cal W}^{s,p}}=\left\|f\right\|_{p}+\left\|f\right\|_{s,p}<\infty\text{,}$
where
$\left\|f\right\|_{s,p}=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\frac{\left|f\left(x\right)-f\left(y\right)\right|^{p}}{\left\|x-y\right\|_{2}^{d+sp}}\text{d}x\text{d}y\text{.}$
It is notable that the fractional Sobolev space ${\cal W}^{s,p}$ is equal to the Besov space ${\cal B}_{p}^{s,p}$ and should be familiar to those working in nonparametric statistics (cf. Leoni,, 2023).
Classes of probability density functions
Let ${\cal P}=\left\{f:\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}\text{ | }\int_{\mathbb{R}^{d}}f\left(x\right)\text{d}x=1\right\}$ denote the set of all probability density functions (PDFs) on $\mathbb{R}^{d}$ Then our target $f_{0}$ is in $\cal P$}. Next, let ${\cal Q}\subset{\cal P}$ be an arbitrary class of PDFs. Then we can define the set of $m$-component mixtures (convex combination of $m$ vertices) of class $\cal Q$} by
$\text{co}_{m}\left({\cal Q}\right)=\left\{f=\sum_{j=1}^{m}\pi_{j}q_{j}\text{ | }q_{j}\in{\cal Q}\text{, }\pi_{j}\geq 0\text{, }\sum_{j=1}^{m}\pi_{j}=1\text{, }j\in\left[m\right]\right\}\text{,}$
The set of all finite mixtures of class $\cal Q$} (all convex combinations) is then written as
$\text{co}\left({\cal Q}\right)=\bigcup_{m\in\mathbb{N}}\text{co}_{m}\left({\cal Q}\right)\text{,}$
where its ${\cal L}_{p}$ closure $\overline{\text{co}\left({\cal Q}\right)}^{p}$ is the set that contains the limit points of all sequences $\left(f_{n}\right)_{n}\subset\text{co}\left({\cal Q}\right)$, with respect to the ${\cal L}_{p}$ topology.
For an arbitrary base PDF $\varphi\in{\cal P}$, we define its dilation (in the language of Cheney & Light,, 2000) by $\nu>0$ by $x\mapsto\varphi_{\nu}=\nu^{d}\varphi\left(\nu x\right)$, and subsequently we define the location–scale class of $\varphi$ by
${\cal P}_{\varphi}=\left\{x\mapsto\varphi_{\nu}\left(x-\mu\right)\text{ | }\mu\in\mathbb{R}^{d}\text{ and }\nu>0\right\}\text{.}$
The location–scale mixtures of $\varphi$ are then $\text{co}\left({\cal P}_{\varphi}\right)$. We will also find use for location-scale classes of $\varphi$, where $\nu$ is fixed, in which case, we shall write ${\cal P}_{\varphi,\nu}=\left\{x\mapsto\varphi_{\nu}\left(x-\mu\right)\text{ | }\mu\in\mathbb{R}^{d}\right\}$.
Convolutions
For two $\lambda$-measurable functions $f,g:\mathbb{R}^{d}\to\mathbb{R}$, we can write their convolution as
$x\mapsto\left(f*g\right)\left(x\right)=\int_{\mathbb{R}^{d}}f\left(y\right)g\left(x-y\right)\text{d}y=\int_{\mathbb{R}^{d}}f\left(x-y\right)g\left(y\right)\text{d}y\text{,}$
where we say that the convolution exists if $\int_{\mathbb{R}^{d}}\left|f\left(y\right)g\left(x-y\right)\right|\text{d}y<\infty$, for $\lambda$-a.e. $x$. In particular, if $f\in{\cal L}_{1}$ and $g\in{\cal L}_{p}$ ($p\geq 1$; e.g., when $f,g\in{\cal P}$), then the convolution exists, $\left\|f*g\right\|_{p}\leq\left\|f\right\|_{1}\left\|g\right\|_{p}$, and $\int_{\mathbb{R}^{d}}\left(f*g\right)\left(x\right)\text{d}x=\int_{\mathbb{R}^{d}}f\left(x\right)\text{d}x\int_{\mathbb{R}^{d}}g\left(y\right)\text{d}y$.
Empirical processes
[…]
3 Approximation rates for location-scale mixtures
3.1 Approximation in Lebesgue spaces
The following theorem, proved in Appendix A, constitutes the main technical result of this manuscript.
Theorem 1.
For $p\in\left(1,\infty\right)$, let $\varphi,f_{0}\in{\cal L}_{p}\cap{\cal P}$ and assume that for some finite $\alpha,K_{1},K_{2}>0$,
$\int_{\mathbb{R}^{d}}\left\|x\right\|_{2}^{\alpha}\varphi\left(x\right)\mathrm{d}x\leq K_{1}\text{,}$ (1)
and
$\left[\int_{\mathbb{R}^{d}}\left|f_{0}\left(x-y\right)-f_{0}\left(x\right)\right|^{p}\mathrm{d}x\right]^{1/p}\leq K_{2}\left\|y\right\|_{2}^{\alpha}\text{.}$ (2)
Then, there exists a sequence $\left(f_{m}\right)_{m}$ with $f_{m}\in\mathrm{co}_{m}\left({\cal P}_{\varphi}\right)$, for each $m\in\mathbb{N}$, such that
$\left\|f_{m}-f_{0}\right\|_{p}\leq\begin{cases}Km^{-\frac{\alpha}{\alpha q+d}}&\text{if }1<p<2\text{,}\\ Km^{-\frac{\alpha q}{2\left(\alpha q+d\right)}}&\text{if }2\leq p<\infty\text{,}\end{cases}$
where $1/p+1/q=1$ and $K$ is a function of $\alpha,K_{1},K_{2},p,d,\varphi$.
3.2 Approximation of functions in Sobolev spaces
Condition (2) is a kind of smoothness assumption enforced on $f_{0}$. For instance, if $f_{0}\in{\cal W}^{1,p}$, it holds with $\alpha=1$ and
$K_{2}=\left[\int_{\mathbb{R}^{d}}\left\|\nabla f_{0}\left(x\right)\right\|_{2}^{p}\text{d}x\right]^{1/p}\text{,}$
and if $f_{0}\in{\cal W}^{s,p}$, then the condition holds with $\alpha=s$ and $K_{2}=\left\|f_{0}\right\|_{s,p}$ (cf. Lemmas 9 and 10).
If we inspect the composition of constant $K$, only $\alpha$ and $K_{2}$ are explicitly dependent on $f_{0}$, with $K_{1}$ being implicitly determined, since $\varphi$ can be chosen conditionally on knowing the smoothness of $f_{0}$ as to satisfy condition (1). In fact, a PDF $\varphi$ with compact support will satisfy (1) for every $\alpha>0$. With these observations in mind, we can state the following concrete result.
Corollary 2.
For $p\in\left(1,\infty\right)$ and $s\in\left(0,1\right]$, if $\varphi\in{\cal L}_{p}\cap{\cal P}$, $f_{0}\in{\cal W}^{s,p}\cap{\cal P}$, and
$\int_{\mathbb{R}^{d}}\left\|x\right\|_{2}^{s}\varphi\left(x\right)\mathrm{d}x\leq K_{1}\text{,}$
then there exists a sequence $\left(f_{m}\right)_{m}$ with $f_{m}\in\mathrm{co}_{m}\left({\cal P}_{\varphi}\right)$, for each $m\in\mathbb{N}$, such that
$\left\|f_{m}-f_{0}\right\|_{p}\leq\begin{cases}Km^{-\frac{s}{sq+d}}&\text{if }1<p<2\text{,}\\ Km^{-\frac{sq}{2\left(sq+d\right)}}&\text{if }2\leq p<\infty\text{,}\end{cases}$
where $K$ is a function of $s,K_{1},p,d,\varphi$ and depends on $f_{0}$ only via $\left\|f_{0}\right\|_{s,p}$ when $s<1$, and $\int_{\mathbb{R}^{d}}\left\|\nabla f_{0}\left(x\right)\right\|_{2}^{p}\mathrm{d}x$ when $s=1$.
As per Theorem 1, the approximation error decreases as a function of $\alpha$ and thus it is desirable to identify concrete sufficient conditions under which $f_{0}$ satisfies (2) with $\alpha>1$, allowing us to guarantee faster approximation rates for such functions. To the best of our knowledge, there are no sufficient conditions such as those of Lemmas 9 and 10 that extend our result. Some closely related conditions are the Calderon–Zygmund kernel assumptions of Propositions 4.1.5 and 4.3.33 of Giné & Nickl, (2016). Unfortunately, for functions with $\alpha>1$, these results require the ability to choose a base function $\varphi$ with zero second moment, which is impossible for any $\varphi\in{\cal P}$.
4 Adaptive least-squares density estimation
Suppose that the random vector $X:\Omega\to\mathbb{R}^{d}$ has density $f_{0}\in{\cal L}_{2}\cap{\cal P}$ in the sense that for each $\mathbb{A}\in\mathfrak{B}\left(\mathbb{R}^{d}\right)$, $\text{P}\left(\omega:X\left(\omega\right)\in\mathbb{A}\right)=\int_{\mathbb{A}}f_{0}\left(x\right)\text{d}x$. If $\left({\cal F}_{n}\right)_{n}$ is a sequence such that ${\cal F}_{n}\subset{\cal L}_{2}\cap{\cal P}$ is a class of PDFs, for each $n\in\mathbb{N}$, and $f_{0}$ is unknown, then we seek to find a sequence of estimators $\left(\hat{f}_{n}\left(X_{1},\dots,X_{n}\right)\right)_{n}$ with each $\hat{f}_{n}\in{\cal F}_{n}$, such that rate of decrease of the expected squared error $\text{E}\left\|\hat{f}_{n}-f_{0}\right\|_{2}^{2}$ can be quantified. For brevity we will write $\hat{f}_{n}=\hat{f}_{n}\left(X_{1},\dots,X_{n}\right)$ for each $n\in\mathbb{N}$, and use similar notation for related estimators. The main result of this section is presented below with the derivation of the result presented in the sequel.
Theorem 3.
For $s\in\left(0,1\right]$, let $\varphi\in\ell^{\infty}\left(\mathbb{R}^{d}\right)\cap{\cal P}$, $f_{0}\in{\cal W}^{s,2}\cap{\cal P}$, $\textsc{vc}\left({\cal P}_{\varphi}\right)<\infty$, and assume that for some finite $K_{1}>0$,
$\int_{\mathbb{R}^{d}}\left\|x\right\|_{2}^{s}\varphi\left(x\right)\mathrm{d}x\leq K_{1}\text{.}$
Then, by taking $m_{n}=\lceil n^{1/2}\rceil$, we can choose a sequence $\left(f_{n}\right)_{n}$, with $\hat{f}_{n}\in\mathrm{co}_{m_{n}}\left({\cal P}_{\varphi}\right)$ depending on $X_{1},\dots,X_{n}$, for each $n\in\mathbb{N}$, such that,
$\mathrm{E}\left\|\hat{f}_{n}-f\right\|_{2}^{2}\lesssim n^{-\frac{s}{2s+d}}\text{.}$
4.1 Proof of Theorem 3
The least-squares criterion
Let ${\cal F}\subset{\cal L}_{2}\cap{\cal P}$. If $f_{0}$ were known, then for each $\epsilon>0$, we can approximate $f_{0}$ by some well-defined $f_{\epsilon}\in{\cal F}$, satisfying
$\left\|f_{\epsilon}-f_{0}\right\|_{2}^{2}\leq\inf_{f\in{\cal F}}\left\|f_{\epsilon}-f_{0}\right\|_{2}^{2}+\epsilon\text{,}$
where $f_{\epsilon}$ is the $\epsilon$-minimizer of the least-squares criterion. We note here that the $\epsilon$-minimizer always exists for each $\epsilon>0$, whereas there may not exist $f^{*}\in{\cal F}$, such that $\left\|f^{*}-f_{0}\right\|_{2}^{2}=\inf_{f\in{\cal F}}\left\|f-f_{0}\right\|_{2}^{2}$.
A decomposition of the least-squares criterion yields
$\left\|f-f_{0}\right\|_{2}^{2}$ $=\int_{\mathbb{R}^{d}}f^{2}\left(x\right)\text{d}x-2\int_{\mathbb{R}^{d}}f\left(x\right)f_{0}\left(x\right)\text{d}x+\int_{\mathbb{R}^{d}}f_{0}^{2}\left(x\right)\text{d}x$
$=\int_{\mathbb{R}^{d}}f^{2}\left(x\right)\text{d}x-2\text{E}f\left(X\right)+\int_{\mathbb{R}^{d}}f_{0}^{2}\left(x\right)\text{d}x\text{,}$
which suggests that a suitable sample estimator of $f_{\epsilon}$ can be obtained by choosing $f_{n,\epsilon}\in{\cal F}$ such that
$\int_{\mathbb{R}^{d}}\hat{f}_{n,\epsilon}^{2}\left(x\right)\text{d}x-2P_{n}\hat{f}_{n,\epsilon}\leq\inf_{f\in{\cal F}}\left[\int_{\mathbb{R}^{d}}f^{2}\left(x\right)\text{d}x-2P_{n}f\right]+\epsilon\text{.}$ (3)
For any $f^{*}\in{\cal F}$, Lemma 11 provides the bound:
$\left\|\hat{f}_{n,\epsilon}-f_{0}\right\|_{2}^{2}\leq\left\|f^{*}-f_{0}\right\|_{2}^{2}+2\left(P_{n}-P\right)\left(\hat{f}_{n,\epsilon}-f^{*}\right)+\epsilon\text{.}$
But
$2\left(P_{n}-P\right)\left(\hat{f}_{n,\epsilon}-f^{*}\right)\leq 4\sup_{f\in{\cal F}}\left|P_{n}f-Pf\right|\text{,}$
leading to the following bound in expectation:
$\text{E}\left\|\hat{f}_{n,\epsilon}-f_{0}\right\|_{2}^{2}\leq\left\|f^{*}-f_{0}\right\|_{2}^{2}+4\text{E}\sup_{f\in{\cal F}}\left|P_{n}f-Pf\right|+\epsilon\text{.}$
Estimators for fixed $m$ and $n$
For each $m,n\in\mathbb{N}$ and $\epsilon>0$, we will consider the construction of an $\epsilon$-minimizer $\hat{f}_{m,n,\epsilon}\in\text{co}_{m}\left({\cal P}_{\varphi,\nu}\right)$, for some optimal $\nu$, such that
$\int_{\mathbb{R}^{d}}\hat{f}_{m,n,\epsilon}^{2}\left(x\right)\text{d}x-2P_{n}\hat{f}_{m,n,\epsilon}$ $\leq\inf_{f\in\text{co}_{m}\left({\cal P}_{\varphi,\nu}\right)}\left[\int_{\mathbb{R}^{d}}f^{2}\left(x\right)\text{d}x-2P_{n}f\right]+\epsilon$
which can be bounded by
$\text{E}\left\|\hat{f}_{m,n,\epsilon}-f_{0}\right\|_{2}^{2}\leq\left\|f_{m,\nu}-f_{0}\right\|_{2}^{2}+4\text{E}\sup_{f\in\text{co}_{m}\left({\cal P}_{\varphi,\nu}\right)}\left|P_{n}f-Pf\right|+\epsilon\text{,}$
with $f_{m,\nu}\in\text{co}_{m}\left({\cal P}_{\varphi,\nu}\right)$, for each $\nu>0$. We recall that an optimal bound of the first term is achieved by setting $\nu$ according to (10):
$\nu_{m}^{*}=\left[\frac{3d\left\|\varphi\right\|_{p}C_{p}}{sqK_{1}K_{2}}\right]^{-\frac{2}{2s+d}}m^{\frac{1}{2s+d}}\text{,}$
yielding the approximation $f_{m,\nu_{m}^{*}}\in\text{co}_{m}\left({\cal P}_{\varphi,\nu_{m}^{*}}\right)$ satisfying
$\left\|f_{m,\nu_{m}^{*}}-f_{0}\right\|_{2}^{2}\leq B_{1}m^{-\frac{2s}{2s+d}}\text{,}$ (4)
for some $B_{1}>0$.
Upon observing that $f\mapsto\left|P_{n}f-Pf\right|$ is a convex function, an application of Lemma 12 implies that
$\text{E}\sup_{f\in\text{co}_{m}\left({\cal P}_{\varphi,\nu}\right)}\left|P_{n}f-Pf\right|=\text{E}\sup_{f\in{\cal P}_{\varphi,\nu}}\left|P_{n}f-Pf\right|\text{.}$ (5)
Further since $\varphi\in\ell^{\infty}\left(\mathbb{R}^{d}\right)$, for each $f\in{\cal P}_{\varphi,\nu}$, there is $\mu\in\mathbb{R}^{d}$ such that $f(x)=\phi_{\nu}(x-\mu)$ which gives:
$\left|f\left(x\right)\right|=\left|\nu^{d}\varphi\left(\nu[x-\mu]\right)\right|\leq\nu^{d}\left\|\varphi\right\|_{\infty}\text{,}$
for every $x\in\mathbb{R}^{d}$. Thus $\bar{F}=\nu^{d}\left\|\varphi\right\|_{\infty}$ is an envelope for ${\cal P}_{\varphi,\nu}$. Then, since $\textsc{vc}\left({\cal P}_{\varphi}\right)<\infty$ and ${\cal P}_{\varphi,\nu}\subset{\cal P}_{\varphi}$, Lemma 13 implies the bound
$\text{E}\sup_{f\in{\cal P}_{\varphi,\nu}}\left|P_{n}f-Pf\right|\leq\nu^{d}\left\|\varphi\right\|_{\infty}\sqrt{\frac{\textsc{vc}\left({\cal P}_{\varphi}\right)}{n}}\text{.}$
Substituting in $\nu_{m}^{*}$ for $\nu$ then implies the bound in $m$ and $n$:
$\text{E}\sup_{f\in\text{co}_{m}\left({\cal P}_{\varphi,\nu}\right)}\left|P_{n}f-Pf\right|\leq\nu_{m}^{*d}\left\|\varphi\right\|_{\infty}\sqrt{\frac{\textsc{vc}\left({\cal P}_{\varphi,\nu}\right)}{n}}\leq\frac{B_{2}}{4}m^{\frac{d}{2s+d}}n^{-1/2}\text{,}$ (6)
for some $B_{2}>0$, achieved by choosing $\epsilon$-minimizers $\hat{f}_{m,n,\epsilon}\in\text{co}_{m}\left({\cal P}_{\varphi,\nu_{m}^{*}}\right)$. Bounds (4) and (6) then combine to obtain
$\text{E}\left\|\hat{f}_{m,n,\epsilon}-f_{0}\right\|_{2}^{2}\leq B_{1}m^{-\frac{2s}{2s+d}}+B_{2}m^{\frac{d}{2s+d}}n^{-1/2}+\epsilon\text{.}$ (7)
Optimizing the rate of $m$
The last step to the proof is to minimize the right-hand side of (7) with respect to $m$, as a function of $n$. Fermat’s condition implies the minimizer in $\mathbb{R}$ is:
$m_{n}^{*}=\frac{2sB_{1}}{dB_{2}}n^{1/2}\text{,}$
Then setting $\hat{f}_{n}=\hat{f}_{\left\lceil m_{n}^{*}\right\rceil,n,\epsilon_{n}}$ and choosing $\epsilon_{n}=B_{3}n^{-\frac{s}{2s+d}}$ for some $B_{3}>0$ yields
$\text{E}\left\|\hat{f}_{n}-f_{0}\right\|_{2}^{2}$ $\leq B_{1}\left\lceil\frac{2sB_{1}}{dB_{2}}n^{1/2}\right\rceil^{-\frac{2s}{2s+d}}+B_{2}\left\lceil\frac{2sB_{1}}{dB_{2}}n^{1/2}\right\rceil^{\frac{d}{2s+d}}n^{-1/2}+B_{3}n^{-\frac{s}{2s+d}}$
$\leq B_{1}\left(\frac{2sB_{1}}{dB_{2}}n^{1/2}\right)^{-\frac{2s}{2s+d}}+B_{2}\left(\frac{4sB_{1}}{dB_{2}}n^{1/2}\right)^{\frac{d}{2s+d}}n^{-1/2}+B_{3}n^{-\frac{s}{2s+d}}$
$=B_{1}\left(\frac{2sB_{1}}{dB_{2}}\right)^{-\frac{2s}{2s+d}}n^{-\frac{s}{2s+d}}+B_{2}\left(\frac{4sB_{1}}{dB_{2}}\right)^{\frac{d}{2s+d}}n^{-\frac{s}{2s+d}}+B_{3}n^{-\frac{s}{2s+d}}$
$=\left[B_{1}\left(\frac{2sB_{1}}{dB_{2}}\right)^{-\frac{2s}{2s+d}}+B_{2}\left(\frac{4sB_{1}}{dB_{2}}\right)^{\frac{d}{2s+d}}+B_{3}\right]n^{-\frac{s}{2s+d}},$
where the second inequality holds whenever $\frac{2sB_{1}}{dB_{2}}n^{1/2}\geq 1$, since $\lceil x\rceil\geq x$ and for $x\geq 1$, $\lceil x\rceil\leq 2x$. Hence, for sufficiently large $n$,
$\text{E}\left\|\hat{f}_{n}-f_{0}\right\|_{2}^{2}\lesssim n^{-\frac{s}{2s+d}}.$ (8)
However, observe that if $m_{n}^{*}$ is replaced by $m_{n}=n^{1/2}$, then by the same argument as above, the estimator $\hat{f}_{n}=\hat{f}_{\left\lceil m_{n}\right\rceil,n,\epsilon_{n}}$ still satisfies (8) but for all $n\geq 1$, as required. That is, for any $n\geq 1$,
$\text{E}\left\|\hat{f}_{n}-f_{0}\right\|_{2}^{2}$ $\leq B_{1}\left\lceil n^{1/2}\right\rceil^{-\frac{2s}{2s+d}}+B_{2}\left\lceil n^{1/2}\right\rceil^{\frac{d}{2s+d}}n^{-1/2}+B_{3}n^{-\frac{s}{2s+d}}$
$\leq B_{1}\left(n^{1/2}\right)^{-\frac{2s}{2s+d}}+B_{2}\left(2n^{1/2}\right)^{\frac{d}{2s+d}}n^{-1/2}+B_{3}n^{-\frac{s}{2s+d}}$
$=B_{1}n^{-\frac{s}{2s+d}}+B_{2}2^{\frac{d}{2s+d}}n^{-\frac{s}{2s+d}}+B_{3}n^{-\frac{s}{2s+d}}$
$=\left[B_{1}+B_{2}2^{\frac{d}{2s+d}}+B_{3}\right]n^{-\frac{s}{2s+d}}.$
5 Discussions
5.1 Vapnik–Chervonenkis classes
We note that the assumption that the VC dimension of the class ${\cal P}_{\varphi}$ be finite in Theorem 3 is not restrictive. In fact, a very large class of functions on $\mathbb{R}^{d}$ can be proved to have finite VC dimension via the fact that if the class of sets ${\cal C}=\left\{\mathbb{G}_{f}\text{ | }f\in\mathcal{F}\right\}$, where $\mathbb{G}_{f}=\left\{\left(x,t\right)\in\mathbb{R}^{d}\times\mathbb{R}\text{ | }t\leq f\left(x\right)\right\}$, is uniformly definable with respect to an o-minimal (order-minimal) structure in the model theoretic sense, then $\textsc{vc}\left({\cal F}\right)<\infty$ (cf., Van den Dries,, 1998, Ch. 5 and Vidyasagar,, 2003, Thm. 10.11).
[…]
Using Vidyasagar, (2003, Thm. 10.11), the fact that $\cal C$} is uniformly definable with respect to an o-minimal structure only permits the qualitative statement that $\textsc{vc}\left({\cal F}\right)<\infty$. Explicit bounds and rates for the VC dimension for o-minimal structres based on the number of operations and number of introduced constants required to compute $f\left(x\right)$ can also be obtained via the results in Anthony & Bartlett, (1999, Ch. 8), Vidyasagar, (2003, Ch. 10), and Montaña & Pardo, (2009).
Based on Vidyasagar, (2003, Thm. 10.11) and the results from Loi, (2010), we can confirm the fact that the location-scale class ${\cal P}_{\varphi}$ based on the Gaussian kernel
$x\mapsto\varphi\left(x\right)=\left(2\pi\right)^{-d/2}\exp\left(-\frac{1}{2}\left\|x\right\|_{2}^{2}\right)$
has finite VC dimension. Other examples include the location-scale classes corresponding to the usual kernel functions in multivariate kernel density estimation, such as the triangular, Epanechnikov, biweight, and triweight kernels (see, e.g., Scott,, 2015, Ch. 6).
5.2 Beyond ${\cal L}_{p}$ spaces
An immediate extension to Theorem 1 is to consider approximation with respect to norms of the form $f\mapsto\left\|f\right\|=\sum_{p\in\mathbb{P}}\left\|f\right\|_{p}$ for some finite set $\mathbb{P}\subset\left(1,\infty\right)$. For such cases, the approximation rate will simply be the slowest rate achieved over all $p\in\mathbb{P}$ from Theorem 1.
Less trivially, we can consider two natural extensions of ${\cal L}_{p}$ spaces. The first extension is to consider the weighted Lebesgue spaces ${\cal L}_{p,w}$, whose norm is determined by a positive function $w:\mathbb{R}^{d}\to\mathbb{R}_{\geq 0}$ via the expression $f\mapsto\left\|f\right\|_{p,w}=\left[\int_{\mathbb{R}^{d}}\left|f\left(x\right)\right|^{p}w\left(x\right)\text{d}x\right]^{1/p}$ for each $p\in\left[1,\infty\right)$, as discussed in Castillo et al., (2016, Sec. 3.8) and appearing often in the approximation theory literature (see, e.g., Mhaskar,, 1996). The required theoretical tools for handling approximation in this setting are the same as those required for the unweighted ${\cal L}_{p}$ spaces (include analogues to Lemmas 68) from which can can infer both qualitative results such as those from Nguyen et al., 2023b as well as the rate results of Theorem 1.
The second extension is to consider variable exponent Lebesgue spaces ${\cal L}_{p\left(\cdot\right)}$ of functions $f$ satisfying $\int_{\mathbb{R}^{d}}\left|f\left(x\right)\right|^{p\left(x\right)}\text{d}x<\infty$, for $p:\mathbb{R}^{d}\to\left[1,\infty\right]$ and whose norm naturally extends the usual ${\cal L}_{p}$ norms when $p\in\left[1,\infty\right]$ is constant (cf. Diening et al.,, 2011 and Cruz-Uribe & Fiorenza,, 2013). Under suitable assumptions, the spaces ${\cal L}_{p\left(\cdot\right)}$ support a Riesz representation theorem and a theory of convolution-based approximation and we can thus extend our results to such spaces with some modifications. Furthermore, the variable exponent Lebesgue spaces naturally define variable exponent Sobolev spaces from which we may derive sufficient conditions for approximation rates such as those of Corollary 2.
Lastly, Riesz representation theorems and convolution-based approximation theorems are known for Sobolev spaces (cf. Adams & Fournier,, 2003 and Leoni,, 2017), from which we can prove qualitative approximation results upon provision of necessary modifications to the proofs of Nguyen et al., 2023b . However, quantitative results such as Theorem 1 may require resolving technicalities that we have not anticipated.
5.3 Faster estimation rates
In the problem setting of estimation of PDFs $f_{0}\in{\cal W}^{s,2}$, for $s>0$, it is well-established that the minimax rate for estimators of the least-squares risk is of order $n^{-\frac{2s}{2s+d}}$ (see, e.g., Klemela,, 2009, Cor. 11.11 and Giné & Nickl,, 2016, Exercise 6.3.2), whose exponent is twice as fast as the rate achieved in Theorem 3. Furthermore, it is also known that this faster rate is achievable, for instance by kernel density or wavelet estimators (cf. Giné & Nickl,, 2016, Sec. 5.1). The question then is whether there is a version of Theorem 3 that achieves the minimax rate.
Following tedious derivations (not included in this version of the text), Massart & Nedelec, (2006, Thm. 2) allows us to prove the following result which can be used in place of Lemma 11.
Lemma 4.
If $f_{0}\in{\cal L}_{\infty}\cap{\cal P}$, ${\cal F}\subset{\cal L}_{2}\cap{\cal L}_{\infty}\cap{\cal P}$ has $\textsc{vc}\left({\cal F}\right)<\infty$ and envelope $\bar{F}$, and $\hat{f}_{n,\epsilon}\in{\cal F}$ satisfies (3), with $\epsilon$ chosen to be an appropriate $\epsilon_{n}$, then
$\mathrm{E}\left\|\hat{f}_{n,\epsilon}-f_{0}\right\|_{2}^{2}\leq 2\inf_{f\in{\cal F}}\left\|f-f_{0}\right\|_{2}^{2}+\left\|\bar{F}\right\|_{\infty}\tilde{O}\left(n^{-1}\right)\text{.}$
If we traverse the proof of Theorem 3 with Lemma 4 in place of Lemma 11, we observe that with only the additional condition $f_{0}\in{\cal L}_{\infty}$, it can be prove that
$\text{E}\left\|\hat{f}_{n}-f_{0}\right\|_{2}^{2}=\tilde{O}\left(n^{-\frac{2s}{2s+d}}\right)\text{,}$
which puts the adaptive least-squares estimator within a logarithmic factor of minimax optimality. Furthermore, when employed in the setting where $f_{0}\in\overline{\text{co}\left({\cal F}\right)}^{2}$, we can show that one can choose estimator sequences $\left(\hat{f}_{n}\right)_{n}\subset\text{co}\left({\cal F}\right)$, such that $\text{E}\left\|\hat{f}_{n}-f_{0}\right\|_{2}^{2}=\tilde{O}\left(n^{-1}\right)$, which is faster than the rate of $O\left(n^{-1/2}\right)$ achieved by Klemelä, (2007) and matches the minimax rate of Kim & Guntuboyina, (2022), up to a logarithmic factor. We currently do not know of any further techniques for improving on these rates.
Appendix
Appendix A Proof of Theorem 1
Let $\varphi_{\nu}$ be a dilate of $\varphi$ and write
$\left\|f_{m}-f_{0}\right\|_{p}\leq\left\|f_{m}-\varphi_{\nu}*f_{0}\right\|_{p}+\left\|\varphi_{\nu}*f_{0}-f_{0}\right\|_{p}\text{.}$ (9)
Then, our proof proceeds in four parts:
  1. 1.
    Prove that for any fixed $\nu>0$, $\varphi_{\nu}*f_{0}\in\overline{\text{co}\left({\cal P}_{\varphi,\nu}\right)}^{p}$ via Lemmas 5 and 6.
  2. 2.
    Use Lemma 7 to construct $(f_{m})$ and then bound $\left\|f_{m}-\varphi_{\nu}*f_{0}\right\|_{p}$ by a function of $\nu$ and $m$.
  3. 3.
    Bound $\left\|\varphi_{\nu}*f_{0}-f_{0}\right\|_{p}$ by a function of $\nu$, using conditions (1) and (2).
  4. 4.
    Combine the bounds and optimize the value of $\nu$ as a function of $m$.
Part 1
[…]
$\int_{\mathbb{R}^{d}}\left(\varphi_{\nu}*f_{0}\right)\left(x\right)h\left(x\right)\text{d}x<\gamma<\int_{\mathbb{R}^{d}}f\left(x\right)h\left(x\right)\text{d}x\text{.}$
But we can write
$\int_{\mathbb{R}^{d}}\left(\varphi_{\nu}*f_{0}\right)\left(x\right)h\left(x\right)\text{d}x$ $=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\varphi_{\nu}\left(x-y\right)f_{0}\left(y\right)\text{d}yh\left(x\right)\text{d}x$
$=\int_{\mathbb{R}^{d}}\int_{\mathbb{R}^{d}}\varphi_{\nu}\left(x-y\right)h\left(x\right)\text{d}xf_{0}\left(y\right)\text{d}y$
$>\int_{\mathbb{R}^{d}}\gamma f_{0}\left(y\right)\text{d}y=\gamma\text{,}$
by Tonelli’s theorem and since $\varphi_{\nu}\left(x-y\right)\in{\cal B}$, which implies the required contradiction.
Part 2
For fixed $\nu>0$, we first observe that by definition for every $g\in\text{co}\left({\cal P}_{\varphi,\nu}\right)$, there exists an $m\in\mathbb{N}$, such that
$g=\sum_{i=1}^{m}\pi_{i}\varphi_{\nu}\left(x-\mu_{i}\right)\text{,}$
for $\mu_{i}\in\mathbb{R}^{d}$ and $\pi_{i}>0$ for each $i\in\left[m\right]$, such that $\sum_{i=1}^{m}\pi_{i}=1$. By Minkowski’s inequality, we have
$\left\|g\right\|_{p}\leq\sum_{i=1}^{m}\pi_{i}\left\|\varphi_{\nu}\left(x-\mu_{i}\right)\right\|_{p}=\left\|\varphi_{\nu}\right\|_{p}\text{.}$
Further, as noted in Section 2, we have that
$\left\|\varphi_{\nu}*f_{0}\right\|_{p}\leq\left\|\varphi_{\nu}\right\|_{p}\left\|f_{0}\right\|_{1}=\left\|\varphi_{\nu}\right\|_{p}\text{,}$
and thus we have the bound
$\left\|g-\varphi_{\nu}*f_{0}\right\|_{p}\leq 2\left\|\varphi_{\nu}\right\|_{p}\text{.}$
But we can write
$\left\|\varphi_{\nu}\right\|_{p}$ $=\left[\int_{\mathbb{R}^{d}}\left[\nu^{d}\varphi\left(\nu x\right)\right]^{p}\text{d}x\right]^{1/p}$
$=\nu^{d\left(p-1\right)/p}\left[\int\varphi^{p}\left(z\right)\text{d}z\right]^{1/p}$
$=\nu^{d/q}\left\|\varphi\right\|_{p}$
by a change in variable and the fact that $q=p/\left(p-1\right)$. Then, since $\varphi_{\nu}*f_{0}\in\overline{\text{co}\left({\cal P}_{\varphi,\nu}\right)}^{p}$, Lemma 7 implies that there exists a sequence $\left(f_{m}\right)_{m}$ with $f_{m}\in\text{co}_{m}\left({\cal P}_{\varphi,\nu}\right)$, for each $m\in\mathbb{N}$, such that
$\left\|f_{m}-\varphi_{\nu}*f_{0}\right\|_{p}\leq\begin{cases}3\nu^{d/q}\left\|\varphi\right\|_{p}C_{p}m^{-1/q}&\text{if }1<p<2\text{,}\\ 3\nu^{d/q}\left\|\varphi\right\|_{p}C_{p}m^{-1/2}&\text{if }2\leq p<\infty\text{,}\end{cases}$
where $C_{p}<\infty$ depends only on $p$, by taking $\epsilon=\nu^{d/q}\left\|\varphi\right\|_{p}$, for simplicity.
Part 3
Start by fixing $\nu>0$ and write:
$\left\|\varphi_{\nu}*f_{0}-f_{0}\right\|_{p}$ $=\left[\int_{\mathbb{R}^{d}}\left|\int_{\mathbb{R}^{d}}\varphi_{\nu}\left(y\right)\left[f_{0}\left(x-y\right)-f_{0}\left(x\right)\right]\text{d}y\right|^{p}\text{d}x\right]^{1/p}$
$\leq\left[\int_{\mathbb{R}^{d}}\left[\int_{\mathbb{R}^{d}}\varphi_{\nu}\left(y\right)\left|f_{0}\left(x-y\right)-f_{0}\left(x\right)\right|\text{d}y\right]^{p}\text{d}x\right]^{1/p}$
$\leq\int_{\mathbb{R}^{d}}\varphi_{\nu}\left(y\right)\left[\int_{\mathbb{R}^{d}}\left|f_{0}\left(x-y\right)-f_{0}\left(x\right)\right|^{p}\text{d}x\right]^{1/p}\text{d}y$
by Jensen’s inequality and Minkowski’s integral inequality Lemma 8, by taking $\varphi_{\nu}\left(y\right)\text{d}y$ as the probability measure on $y$. Then, by (2), we have
$\int_{\mathbb{R}^{d}}\varphi_{\nu}\left(y\right)\left[\int_{\mathbb{R}^{d}}\left|f_{0}\left(x-y\right)-f_{0}\left(x\right)\right|^{p}\text{d}x\right]^{1/p}\text{d}y$ $\leq K_{2}\int_{\mathbb{R}^{d}}\left\|y\right\|_{2}^{\alpha}\varphi_{\nu}\left(y\right)\text{d}y$
$=K_{2}\int_{\mathbb{R}^{d}}\left\|y\right\|_{2}^{\alpha}\nu^{d}\varphi\left(\nu y\right)\text{d}y$
$=\frac{K_{2}}{\nu^{\alpha}}\int_{\mathbb{R}^{d}}\left\|z\right\|_{2}^{\alpha}\varphi\left(z\right)\text{d}z\leq K_{1}K_{2}\nu^{-\alpha}\text{,}$
using condition (1) and a change in variable.
Part 4
In the case when $p\in\left(1,2\right)$, we have from (9) that
$\left\|f_{m}-f_{0}\right\|_{p}\leq 3\nu^{d/q}\left\|\varphi\right\|_{p}C_{p}m^{-1/q}+K_{1}K_{2}\nu^{-\alpha}\text{.}$
By minimizing the right-hand side with respect to $\nu$, we have the minimizer
$\nu_{m}^{*}=\left[\frac{3d\left\|\varphi\right\|_{p}C_{p}}{\alpha qK_{1}K_{2}}\right]^{-\frac{q}{\alpha q+d}}m^{\frac{1}{\alpha q+d}}\text{,}$
which implies the optimal bound $Km^{-\frac{\alpha}{\alpha q+d}}$, where
$K=3\left\|\varphi\right\|_{p}C_{p}\left[\frac{3d\left\|\varphi\right\|_{p}C_{p}}{\alpha qK_{1}K_{2}}\right]^{-\frac{d}{\alpha q+d}}+K_{1}K_{2}\left[\frac{3d\left\|\varphi\right\|_{p}C_{p}}{\alpha qK_{1}K_{2}}\right]^{\frac{\alpha q}{\alpha q+d}}\text{.}$
Similarly, when $p\in\left[2,\infty\right)$, we have the bound
$\left\|f_{m}-f_{0}\right\|_{p}\leq 3\nu^{d/q}\left\|\varphi\right\|_{p}C_{p}m^{-1/2}+K_{1}K_{2}\nu^{-\alpha}\text{,}$
which is minimized at
$\nu^{*}=\left[\frac{3d\left\|\varphi\right\|_{p}C_{p}}{\alpha qK_{1}K_{2}}\right]^{-\frac{q}{\alpha q+d}}m^{\frac{q}{2\left(\alpha q+d\right)}}\text{.}$ (10)
This yields the optimal bound $Km^{-\frac{\alpha q}{2\left(\alpha q+d\right)}}$.
Appendix B External technical results
Lemma 5 (Geometric Hahn–Banach theorem).
Suppose that $\mathbb{X}$ is a real Banach space and $\mathbb{A},\mathbb{B}\subset\mathbb{X}$ are non-empty, disjoint, and convex sets. Further assume that $\mathbb{A}$ is compact and $\mathbb{B}$ is closed. Then, there exists a $h\in\mathbb{X}^{*}$ (the dual of $\mathbb{X}$) and $\gamma\in\mathbb{R}$, such that for every $a\in\mathbb{A}$ and $b\in\mathbb{B}$,
$h\left(a\right)<\gamma<h\left(b\right)\text{.}$
Lemma 6 (Riesz representation theorem).
For $p\in\left(1,\infty\right)$ and $1/p+1/q=1$, if $h\in\left({\cal L}^{p}\right)^{*}$, then there exists a unique $g\in{\cal L}^{q}$ such that
$h\left(f\right)=\int_{\mathbb{R}^{d}}f\left(x\right)g\left(x\right)\text{d}x\text{,}$
for every $f\in{\cal L}^{p}$.
Lemma 7.
Let ${\cal A}\subset{\cal L}^{p}$ for $p\in\left(1,\infty\right)$ and suppose that $f\in\overline{\mathrm{co}\left({\cal A}\right)}^{p}$. Further, assume that there exists a $B>0$ such that for all $g\in{\cal A}$, $\left\|g-f\right\|_{p}\leq B$. Then, for each $\epsilon>0$, there exists a sequence $\left(g_{m}\right)_{m}$ with $g_{m}\in\mathrm{co}_{m}\left({\cal A}\right)$, for each $m\in\mathbb{N}$, such that
$\left\|g_{m}-f\right\|_{p}\leq\begin{cases}\left(B+\epsilon\right)C_{p}m^{-1/q}&\text{if }1\leq p<2\text{,}\\ \left(B+\epsilon\right)C_{p}m^{-1/2}&\text{if }2\leq p<\infty\text{,}\end{cases}$
where $C_{p}<\infty$ is a fixed constant depending only on $p$ and $1/p+1/q=1$.
Lemma 8 (Minkowski’s integral inequality).
Let $\left(\mathbb{X},\mathfrak{X},\mu_{1}\right)$ and $\left(\mathbb{Y},\mathfrak{Y},\mu_{2}\right)$ be $\sigma$-finite measure spaces, and let $f:\mathbb{X}\times\mathbb{T}\to\bar{\mathbb{R}}$ be $\mathfrak{X}\otimes\mathfrak{Y}$-measurable. Then, for each $p\in\left[1,\infty\right)$,
$\left[\int_{\mathbb{X}}\left[\int_{\mathbb{Y}}\left|f\left(x,y\right)\right|\mu_{2}\left(\mathrm{d}y\right)\right]^{p}\mu_{1}\left(\mathrm{d}x\right)\right]^{1/p}\leq\int_{\mathbb{Y}}\left[\int_{\mathbb{X}}\left|f\left(x,y\right)\right|^{p}\mu_{1}\left(\mathrm{d}x\right)\right]^{1/p}\mu_{2}\left(\mathrm{d}x\right)\text{.}$
Lemma 9.
Let $p\in\left[1,\infty\right)$ and $f\in{\cal W}^{1,p}$. Then, for all $y\in\mathbb{R}^{d}$,
$\left[\int_{\mathbb{R}^{d}}\left|f\left(x-y\right)-f\left(x\right)\right|^{p}\mathrm{d}x\right]^{1/p}\leq\left[\int_{\mathbb{R}^{d}}\left\|\nabla f\left(x\right)\right\|_{2}^{p}\text{d}x\right]^{1/p}\left\|y\right\|_{2}\text{.}$
Lemma 10.
Let $p\in\left[1,\infty\right)$ and $s\in\left(0,1\right)$. Then, for any $y\in\mathbb{R}^{d}$,
$\left[\int_{\mathbb{R}^{d}}\left|f\left(x-y\right)-f\left(x\right)\right|^{p}\mathrm{d}x\right]^{1/p}\leq\left\|f\right\|_{s,p}\left\|y\right\|_{2}^{s}\text{.}$
Lemma 11.
If ${\cal F}\subset{\cal L}_{2}\cap{\cal P}$ and $\hat{f}_{n,\epsilon}\in{\cal F}$ satisfies (3), then for each $f^{*}\in{\cal F}$ and $\epsilon>0$,
$\left\|\hat{f}_{n,\epsilon}-f_{0}\right\|_{2}^{2}\leq\left\|f^{*}-f_{0}\right\|_{2}^{2}+2\left(P_{n}-P\right)\left(\hat{f}_{n,\epsilon}-f^{*}\right)+\epsilon\text{.}$
Lemma 12.
Let $f:\mathbb{X}\to\mathbb{R}$ be a convex function and
$\mathbb{X}=\left\{\sum_{j=1}^{m}\pi_{j}a_{j}\,|\,a_{j}\in\mathbb{A}\text{, }\pi_{j}\geq 0\text{, }j\in\left[m\right]\text{, }m\in\mathbb{N}\right\}$
be the convex combinations of some set $\mathbb{A}$. Then,
$\sup_{x\in\mathbb{X}}f\left(x\right)=\sup_{x\in\mathbb{A}}f\left(x\right)\text{,}$
where the left-hand side is attained only when the right-hand side is attained.
Lemma 13.
If ${\cal F}\subset\ell^{\infty}\left(\mathbb{R}^{d}\right)$ has an envelope $\bar{F}$ and satisfies $\textsc{vc}\left({\cal F}\right)<\infty$, then
$\mathrm{E}\sup_{f\in{\cal F}}\left|P_{n}f-Pf\right|\lesssim\sqrt{\frac{\textsc{vc}\left({\cal F}\right)}{n}}\left\|\bar{F}\right\|_{P,2}\text{.}$
Appendix C Sources of external results
We provide some references to sources of the external results which may have been paraphrased for the context of this work. Lemma 5 appears as Theorem 21.2 in Robinson, (2020). Lemma 6 appears as Corollary 21.6 in Schilling, (2017). Lemma 7 appears as Corollary 3.6 in Donahue et al., (1997). Lemma 8 appears as Theorem 14.16 in Schilling, (2017). Lemma 9 appears as Lemma 12.19 in Leoni, (2017). Lemma 10 appears as Lemma 6.14 in Leoni, (2023). Lemma 13 is stated in Section 2.14.4 of van der Vaart & Wellner, (2023).